# Fractal counterparts of density combinatorics theorems in continued fractions

### v1

@@ 2.2. Lower bounds of Hausdorff dimension
  • $E_{n-1}$ is the union of finitely many closed intervals of positive length, called $(n-1)$-th level intervals;
  • for each $(n-1)$-th level interval $I$, the number of $n$-th intervals contained in $E_{n}\cap I$ is at least $r_{n}$, and all of these $n$-th intervals are separated by gaps of length at least $\delta_{n}$, where $r_{n}\geq 2$ and $0<\delta_{n+1}<\delta_{n}$.
We call $F$ a Moran fractal with profile $(r_{n})_{n=1}^{\infty}$, $(\delta_{n})_{n=1}^{\infty}$. The Hausdorff dimension of a Moral fractal is estimated from below with its profile by the next lemma.
Lemma 2.2 ([8] pp.71–72).
Let $F\subset[0,1]$ be a Moran fractal with profile $(r_{n})_{n=1}^{\infty}$, $(\delta_{n})_{n=1}^{\infty}$. Then
@@
Lemma 2.3 ([8] Proposition 3.3).
Let $F\subset[0,1]$ and let $f\colon F\to[0,1]$ be Hölder continuous with exponent $\gamma\in(0,1]$. Then $\dim_{\rm H}F\geq\gamma\cdot\dim_{\rm H}f(F).$
@@ 2.3. Convergence exponent
We call $\tau(S)$ a convergence exponent of $S$. Note [^2] The convergence exponent was originally defined for infinite sequences of positive reals, see [30, §$\S$2]. Then the value ranges from 00 to $\infty$. that $\tau(S)\in[0,1]$. If we write $S=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots$ then
@@
see [30, §$\S$2]. The convergence exponent is often used in fractal dimension theory, see [15, 40] for example. The next lemma can be proved by a standard covering argument by fundamental intervals.
Lemma 2.5 ([15] Corollary 1).
@@
The inequality in Lemma 2.5 is actually the equality for any infinite subset $S$ of $\mathbb{N}$ (see [40]), but we do not need this fact. The equality for $S=\mathbb{N}$ is well-known as Good’s theorem [13].
The next lemma links the convergence exponent to relative upper density. The case $A\neq S$ will only be used in the proof of Theorem 1.13.
@@
If $S\subset\mathbb{N}$ has polynomial density with exponent $\alpha$, then for any $A\subset S$ with $\overline{d}(A|S)>0$ we have
@@
In the case $A=S$, the same argument as above yields the reverse inequality $\tau(A)\geq 1/\alpha.$ In the case $A\neq S$, the same argument does not work since the positivity of $\overline{d}(A|S)$ does not imply that of $\limsup_{n\to\infty}\#(A\cap[1,a_{n}])/\#(S\cap[1,a_{n}])$. Hence we argue as follows. Take a strictly increasing sequence $(N_{k})_{k=1}^{\infty}$ of positive integers such that
@@
Let $\varepsilon\in(0,1/\alpha)$. Since $S$ has polynomial density with exponent $\alpha$, there exist $C_{1}>0$ and $\beta\geq 0$ such for all sufficiently large $k$ we have
#$\begin{split}\frac{\#(A\cap[N_{k}^{\varepsilon},N_{k}])}{\#(S\cap[1,N_{k}])}&\geq\frac{\#(A\cap[1,N_{k}])}{\#(S\cap[1,N_{k}])}-\frac{\#(A\cap[1,N_{k}^{\varepsilon}])}{\#(S\cap[1,N_{k}])}\\ &\geq\frac{\overline{d}(A|S)}{2}-\frac{(\log N_{k})^{\beta}}{C_{1}N_{k}^{1/\alpha-\varepsilon}}\geq\frac{\overline{d}(A|S)}{3}.\end{split}$
@@
#$\#(A\cap[N_{k}^{\varepsilon},N_{k}])\geq\frac{\overline{d}(A|S)}{3}\#(S\cap[1,N_{k}])\geq\frac{\overline{d}(A|S)}{3}\frac{C_{1}N_{k}^{1/\alpha}}{(\log N_{k})^{\beta}}.$
@@
The next result obtained in the multifractal analysis of convergence exponent for continued fractions [9] is a key ingredient in the proof of Theorem 1.13.
Lemma 2.8 ([9], Theorem 2.2).
We have
[…]
3. Key constructions
This section is dedicated to key constructions in the proofs of the main results. In §$\S$3.1 we introduce and construct an extreme seed set. In §$\S$3.2 we describe a procedure of constructing a new Moran-like fractal of $(0,1)\setminus\mathbb{Q}$ by adding digits into a given seed set. In §$\S$3.3 we prove that an elimination map associated with the construction in §$\S$3.2 is almost Lipschitz.
3.1. Construction of extreme seed sets
Let $S\subset\mathbb{N}$ have polynomial density. For an infinite subset $S_{*}$ of $S$ and positive reals $t,L$ satisfying $t>L>1$, define
$R_{t,L}(S_{*})=\{x\in E\colon a_{n}(x)\in S_{*}\cap[t^{n},Lt^{n}]\ \text{ for every}\ n\geq 1\}.$
Non-empty sets of this form are called seed sets. Seed sets are Moran-like fractals, and their Hausdorff dimensions do not exceed $1/(2\alpha)$ by Lemmas 2.5 and 2.7. In fact, one can decrease the relative upper density of $S_{*}$ without decreasing the Hausdorff dimension of the corresponding seed set. We say the seed set $R_{t,L}(S_{*})$ is an extreme seed set if $S_{*}$ does not contain consecutive integers, and
$\overline{d}(S_{*}|S)=0\ \text{ and }\ \dim_{\rm H}R_{t,L}(S_{*})=\frac{1}{2\alpha}.$
Extreme seed sets are of maximal Hausdorff dimension and least relative density of digits. The assumption of not containing consecutive integers is not essential, but a little facilitates the proof of Proposition 3.5 in §$\S$3.2.
@@ 3.1. Construction of extreme seed sets
If $S\subset\mathbb{N}$ has polynomial density, then there exist $S_{*}\subset S$ and positive reals $t,L$ such that $t>L>1$, $t>2$ and $R_{t,L}(S_{*})$ is an extreme seed set.
For a proof of this proposition we need two preliminary lemmas. First, define a function $\nu\colon[1,\infty)\to\mathbb{N}$ by
@@
This is a slow increasing function that has the following upper bound.
@@
We claim that $\nu(t^{n})\leq n\log t$ for every $n\in\mathbb{N}$ but finitely many ones. For otherwise the definition of $\nu$ in (3.1) would yield $t^{n}>\lfloor n\log t\rfloor!$ for infinitely many $n$, which is a contradiction. There exists a constant $c(t)\geq 1$ such that $\nu(t^{n})\leq c(t)n\log t$ for every $n\in\mathbb{N}$. Take $\lambda>1$ such that $c(t)\leq(\log t)^{\lambda-1}$. ∎
The next general lemma allows us to extract from an arbitrary infinite subset of $\mathbb{N}$ a set of zero relative density with definite growth bounds.
Lemma 3.3.
$\frac{1}{2\nu(\Delta)}\leq\frac{\#\{1\leq n\leq\Delta\colon a_{n}\in S_{*}\}}{\Delta}\leq\frac{3}{\nu(\Delta)}$ (3.2)
for all sufficiently large $\Delta\geq 1$. In particular, $\overline{d}(S_{*}|S)=0$.
Proof.
@@
The following interpretation of $Q$ should be useful along the way. Split $[2,\infty)=\bigcup_{k=2}^{\infty}[k!,(k+1)!)$. For each integer $k\geq 2$, further split $[k!,(k+1)!)$ into consecutive left-closed, right-open subintervals containing exactly $k$ integers:
$[k!,(k+1)!)=[k!,k!+k)\cup[k!+k,k!+2k)\cup\cdots\cup[k!+(k!-1)k,(k+1)!).$
@@
For every $\Delta\geq 1$ we have
#$\#\{1\leq n\leq\Delta\colon a_{n}\in S_{*}\}=\#(Q\cap[1,\Delta]).$
If $\Delta\geq 1$ is sufficiently then we have
#$\begin{split}\#(Q\cap[1,\Delta])&=\#(Q\cap[1,(k-1)!))+\#(Q\cap[(k-1)!,k!))+\#(Q\cap[k!,\Delta])\\ &\leq\#(Q\cap[1,(k-1)!))+\frac{\#([(k-1)!,k!)\cap\mathbb{N})}{k-1}+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}\\ &\leq(k-1)!+(k-1)!+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}\leq\frac{3\Delta}{\nu(\Delta)},\end{split}$
@@
#$\begin{split}\#(Q\cap[1,\Delta])&\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{i}+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}-1\\ &\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{k}+\frac{\#([k!,\Delta]\cap\mathbb{N})}{k}-1\\ &=\frac{\#([2,\Delta]\cap\mathbb{N})}{k}-1\geq\frac{\Delta}{2\nu(\Delta)},\end{split}$
where $k\geq 1$ is the integer with $\Delta\in[k!,(k+1)!)$. This verifies (3.2).
The ordering of elements of $S$ implies #$\#(S_{*}\cap[1,N])=\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}$ for every $N\geq 1$ with $S\cap[1,N]\neq\emptyset$. By the second inequality in (3.2), for all sufficiently large $N\geq 1$ we have
#$\frac{\#(S_{*}\cap[1,N])}{\#(S\cap[1,N])}=\frac{\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}}{\#(S\cap[1,N])}\leq\frac{3}{\nu(\#(S\cap[1,N]))}.$
@@
Proof of Proposition 3.1.
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha$ and write $S=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots.$ Let $S_{*}$ be a subset of $S$ for which the conclusion of Lemma 3.3 holds. It suffices to show that $\dim_{\rm H}R_{t,L}(S_{*})\geq 1/(2\alpha)$ for some $t,L$ with $t>L>1$ and $t>2$.
By the ordering of elements of $S$, for any $N\geq 1$ with $S\cap[1,N]\neq\emptyset$ we have #$\#(S_{*}\cap[1,N])=\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}$. Moreover (3.2) gives
#$\begin{split}\frac{\#(S\cap[1,N])}{2\nu(\#(S\cap[1,N]))}\leq\#(S_{*}\cap[1,N])&\leq\frac{3\#(S\cap[1,N])}{\nu(\#(S\cap[1,N]))}.\end{split}$ (3.3)
@@
Using (3.3), (3.5) and then (3.6), (3.7), for each integer $n\geq 1$ we have
#$\begin{split}\#(S_{*}\cap[t^{n},Lt^{n}])&\geq\#(S_{*}\cap[1,Lt^{n}])-\#(S_{*}\cap[1,t^{n}])\\ &\geq\frac{\#(S\cap[1,Lt^{n}])}{2\nu(\#(S\cap[1,Lt^{n}]))}-\frac{3\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}\\ &\geq\frac{C_{1}C_{2}^{-1}2^{-\beta}L^{1/\alpha}\#(S\cap[1,t^{n}])}{2(\nu(\#(S\cap[1,t^{n}])+1)}-\frac{3\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}.\end{split}$
@@
#$\#(S\cap[1,t^{n}])\left(\frac{C_{1}C_{2}^{-1}2^{-\beta}L^{1/\alpha}}{3\nu(\#(S\cap[1,t^{n}]))}-\frac{3}{\nu(\#(S\cap[1,t^{n}]))}\right),$
which equals #$\#(S\cap[1,t^{n}])/\nu(\#(S\cap[1,t^{n}]))$ by (3.4). Set $\rho=\beta+\lambda$. Using (3.8) and the assumption that $S$ has polynomial density, we obtain
#$\begin{split}\#(S_{*}\cap[t^{n},Lt^{n}])\geq\frac{\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}&\geq C_{1}\frac{t^{\frac{n}{\alpha}}}{(n\log t)^{\beta}}\cdot\frac{1}{C_{1}(n\log t)^{\lambda}}\\ &=\frac{t^{\frac{n}{\alpha}}}{(n\log t)^{\rho}}.\end{split}$ (3.9)
@@
where the bar denotes the closure. Since $t>L$, the intervals $[t^{n},Lt^{n}]$, $n\in\mathbb{N}$ are pairwise disjoint. The Moran fractal $F$ is contained in $R_{t,L}(S_{*})$ up to countable sets consisting of the boundary points of the level intervals. For every $n\geq 2$ and every $(n-1)$-th level interval of $F$, the number of $n$-th level intervals contained in that interval is at least ${\lfloor t^{\frac{n}{\alpha}}/(n\log t)^{\rho}\rfloor}$ by (3.9). Since $S_{*}$ does not contain consecutive integers, these $n$-th level intervals are separated at least by some $n$-th fundamental intervals. Hence, the gap between the $n$-th level intervals is at least
$\min\{|I(a_{1},\ldots,a_{n})|\colon t^{i}\leq a_{i}\leq Lt^{i},1\leq i\leq n\}.$
By Lemma 2.1 and $t>L$, we have
@@
So, $F$ is a Moran fractal with profile $({\lfloor t^{\frac{n}{\alpha}}/(n\log t)^{\rho}\rfloor})_{n=1}^{\infty}$, $((1/2)\prod_{i=1}^{n}(Lt^{i+1})^{-2})_{n=1}^{\infty}$. Lemma 2.2 gives
@@
Let us agree that for two eventually positive functions $f$, $g$ on $\mathbb{N}$, the expression $f(n)\sim g(n)$ indicates that $f(n)/g(n)\to 1$ as $n\to\infty$. Direct calculations show
$\log\left(2\prod_{i=1}^{n-1}{\left\lfloor\frac{t^{\frac{i}{\alpha}}}{(i\log t)^{\rho}}\right\rfloor}\right)$ $=\log 2+\sum_{i=1}^{n-1}\log\left\lfloor\frac{t^{\frac{i}{\alpha}}}{(i\log t)^{\rho}}\right\rfloor$
$\sim\log 2+\sum_{i=1}^{n-1}\left(\log t^{\frac{i}{\alpha}}-\log(i\log t)^{\rho}\right)\sim\frac{1}{2\alpha}n^{2}\log t,$
@@
Corollary 3.4.
If $S\subset\mathbb{N}$ has polynomial density with exponent $\alpha$, then
$\dim_{\rm H}\{x\in J\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}=\dim_{\rm H}\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}=\frac{1}{2\alpha}.$
@@
3.2. Adding digits into seed sets

### v2

@@ 2.2. Lower bounds of Hausdorff dimension
  • $E_{n-1}$ is the union of finitely many closed intervals in $[0,1]$ of positive diameter, called (n1)(n-1)-th level intervals;
  • for each (n1)(n-1)-th level interval $I$, the number of $n$-th intervals contained in $E_{n}\cap I$ is at least $r_{n}$, and all of these $n$-th intervals are separated by gaps of diameter at least $\delta_{n}$, where $r_{n}\geq 2$ and $0<\delta_{n+1}<\delta_{n}$.
We call $F$ a Moran fractal with parameters $(r_{n})_{n=1}^{\infty}$, $(\delta_{n})_{n=1}^{\infty}$. The Hausdorff dimension of a Moral fractal is estimated from below with its parameters by the next lemma.
Lemma 2.2 ([8, pp.71–72]).
Let $F\subset[0,1]$ be a Moran fractal with parameters $(r_{n})_{n=1}^{\infty}$, $(\delta_{n})_{n=1}^{\infty}$. Then
@@
Lemma 2.3 ([8, Proposition 3.3]).
Let $F\subset[0,1]$ and let $f\colon F\to[0,1]$ be Hölder continuous with exponent $\gamma\in(0,1]$. Then
$\dim_{\rm H}F\geq\gamma\cdot\dim_{\rm H}f(F).$
@@ 2.3. Convergence exponent
We call $\tau(S)$ a convergence exponent of $S$. Note that $\tau(S)\in[0,1]$. If we write $S=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots$ then
@@
see [34, Section 2]. The convergence exponent is often used in fractal dimension theory, see [16, 45] for example. The next lemma can be proved by a standard covering argument by fundamental intervals.
Lemma 2.5 ([16, Corollary 1]).
@@
The inequality in Lemma 2.5 is actually the equality for any infinite subset $S$ of $\mathbb{N}$, known as a solution of Hirst’s conjecture [45]. The inequality suffices for the proofs of our main results. The equality for $S=\mathbb{N}$ is known as Good’s theorem [14, Theorem 1].
The next lemma links the convergence exponent to relative upper density. We will mostly use it with $A=S$ in combination with Lemma 2.5, in order to obtain an upper bound of Hausdorff dimension. The case $A\subsetneq S$ will only be used in the proof of Theorem 1.13.
@@
If $S\subset\mathbb{N}$ has polynomial density with exponent $\alpha\geq 1$, then for any $A\subset S$ with $\overline{d}(A|S)>0$ we have
@@
In the case $A=S$, the same argument as above yields the reverse inequality $\tau(A)\geq 1/\alpha.$ In the case $A\subsetneq S$, the same argument does not work since $\overline{d}(A|S)>0$ does not imply $\limsup_{n\to\infty}\#(A\cap[1,a_{n}])/\#(S\cap[1,a_{n}])>0$. Take a strictly increasing sequence $(N_{k})_{k=1}^{\infty}$ of positive integers such that
@@
Let $\varepsilon\in(0,1/\alpha)$. Since $S$ has polynomial density with exponent $\alpha\geq 1$, there exist $C_{1}>0$ and $\beta\geq 0$ such that for all sufficiently large $k$ we have
$\begin{split}\frac{\#(A\cap[N_{k}^{\varepsilon},N_{k}])}{\#(S\cap[1,N_{k}])}&\geq\frac{\#(A\cap[1,N_{k}])}{\#(S\cap[1,N_{k}])}-\frac{\#(A\cap[1,N_{k}^{\varepsilon}])}{\#(S\cap[1,N_{k}])}\\ &\geq\frac{\overline{d}(A|S)}{2}-\frac{(\log N_{k})^{\beta}}{C_{1}N_{k}^{1/\alpha-\varepsilon}}\geq\frac{\overline{d}(A|S)}{3}.\end{split}$
@@
$(A\cap[N_{k}^{\varepsilon},N_{k}])\geq\frac{\overline{d}(A|S)}{3}\#(S\cap[1,N_{k}])\geq\frac{\overline{d}(A|S)}{3}\frac{C_{1}N_{k}^{1/\alpha}}{(\log N_{k})^{\beta}}.$
@@
Remark 2.8.
The difference between the upper density and the upper Banach density appear in their relations with convergence exponent. By Lemma 2.7, if $A\subset\mathbb{N}$ and $\overline{d}(A)>0$ then $\tau(A)=1$. It is easy to see that $d_{\rm B}(A)>0$ does not imply $\tau(A)=1$: for example, for the set $A=\{\left(\bigcup_{n\in\mathbb{N}}[n^{2},n^{2}+\sqrt{n}]\right)\cap\mathbb{N}\}$ we have $d_{\rm B}(A)=1$ and $\tau(A)=3/4$.
The next result on the multifractal analysis of convergence exponent for continued fractions [9] is a key ingredient in the proof of Theorem 1.13.
Lemma 2.9 ([9, (2.2) on p.1899]).
For all $0\leq c\leq 1$ we have
$\dim_{\rm H}\{x\in J\colon\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})=c\}=\frac{1-c}{2}.$
Remark 2.10.
The difference between a fractal structure of $J$ and that of $E$ appear in their relations with the multifractal spectrum of convergence exponent. Taking $S=\mathbb{N}$ in (1.3) and combining the result with Lemma 2.7, we get
$\dim_{\rm H}\{x\in E\colon\tau(\{a_{n}(x)\colon n\in\mathbb{N}\})=1\}=\frac{1}{2},$
which is a sharp contrast to the equality in Lemma 2.9 with $c=1$.
3. Three-step general construction
In this section we perform the three-step general construction of Moran fractals in $E$ described in Section 1.3. In Section 3.1 we introduce extreme seed sets that we are going to construct in Step 2. In Section 3.2 we complete Step 1, and in Section 3.3 complete Step 2. The remaining two subsections are on Step 3 that is most technical. In Section 3.4, we describe a procedure of constructing a new Moran fractal in $E$ by adding digits into a given seed set. In Section 3.5 we prove that an elimination map associated with the construction in Section 3.4 is almost Lipschitz.
3.1. Extreme seed sets
Let $S$ be an infinite subset of $\mathbb{N}$. For an infinite subset $K$ of $S$ and positive reals $t,L$ satisfying $t>L>1$, define
$R_{t,L}(K)=\{x\in E\colon a_{n}(x)\in K\cap[t^{n},Lt^{n}]\ \text{ for every}\ n\geq 1\}.$ (3.1)
Note that $R_{t,L}(K)\subset J$. Non-empty sets of this form are called seed sets. We say the seed set $R_{t,L}(K)$ is an extreme seed set associated with $S$ if the following conditions hold:
  • (A1)
    $S$ has polynomial density with exponent $\alpha\geq 1$;
  • (A2)
    $K$ does not contain consecutive integers;
  • (A3)
    $\overline{d}(K|S)=0;$
  • (A4)
    $\dim_{\rm H}R_{t,L}(K)=1/(2\alpha)$.
By Lemmas 2.5 and 2.7, if (A1) holds then the Hausdorff dimension of any seed set associated with $S$ does not exceed $1/(2\alpha)$. Hence, extreme seed sets are of least relative density of digits (A3) and of maximal Hausdorff dimension (A4). Condition (A1) is not essential, but a little facilitates the proof of Proposition 3.5 in Section 3.4.
@@ 3.1. Extreme seed sets
If $S\subset\mathbb{N}$ has polynomial density, then there exist $S_{*}\subset S$ and positive reals $t,L$ such that $t>L>1$, $t>2$ and $R_{t,L}(S_{*})$ is an extreme seed set associated with $S$.
Prior to a proof of Proposition 3.1, as Step 1 we construct in Section 3.2 a relatively thin subset of a given infinite subset of $\mathbb{N}$. The proof of Proposition 3.1 is given in Section 3.3 as Step 2.
3.2. Construction of a relatively thin subset (Step 1)
The next general lemma allows us to extract from an arbitrary infinite subset of $\mathbb{N}$ a set of zero relative density with definite growth bounds. Define a function $\nu\colon[1,\infty)\to\mathbb{N}$ by
@@ 3.2. Construction of a relatively thin subset (Step 1)
Lemma 3.2.
$\frac{1}{2\nu(\xi)}\leq\frac{\#\{1\leq n\leq\xi\colon a_{n}\in S_{*}\}}{\xi}\leq\frac{3}{\nu(\xi)}$ (3.3)
for all sufficiently large $\xi\geq 1$. In particular, $\overline{d}(S_{*}|S)=0$.
For a proof of Lemma 3.2, we first show that $\nu$ is a slow increasing function with the following upper bound.
@@
We claim that $\nu(t^{n})\leq n\log t$ for all $n\in\mathbb{N}$ but finitely many ones. For otherwise the definition of $\nu$ in (3.2) would yield $t^{n}>\lfloor n\log t\rfloor!$ for infinitely many $n$, which is a contradiction. There exists a constant $c(t)\geq 1$ such that $\nu(t^{n})\leq c(t)n\log t$ for every $n\in\mathbb{N}$. Take $\lambda>1$ such that $c(t)\leq(\log t)^{\lambda-1}$. ∎
Proof of Lemma 3.2.
@@
The following interpretation of $Q$ should be useful along the way. Split $2,\infty)=\bigcup_{k=2}^{\infty}[k!,(k+1)!)$. For each integer $k\geq 2$, further split $[k!,(k+1)!)$ into consecutive left-closed, right-open subintervals containing exactly $k$ integers:
$k!,(k+1)!)=[k!,k!+k)\cup[k!+k,k!+2k)\cup\cdots\cup[k!+(k!-1)k,(k+1)!).$
@@
For every $\xi\geq 1$ we have
$\{1\leq n\leq\xi\colon a_{n}\in S_{*}\}=\#(Q\cap[1,\xi]).$
If $\xi\geq 1$ is sufficiently large then we have
$\begin{split}\#(Q\cap[1,\xi])&=\#(Q\cap[1,(k-1)!))+\#(Q\cap[(k-1)!,k!))+\#(Q\cap[k!,\xi])\\ &\leq\#(Q\cap[1,(k-1)!))+\frac{\#([(k-1)!,k!)\cap\mathbb{N})}{k-1}+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}\\ &\leq(k-1)!+(k-1)!+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}\leq\frac{3\xi}{\nu(\xi)},\end{split}$
@@
$\begin{split}\#(Q\cap[1,\xi])&\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{i}+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}-1\\ &\geq\sum_{i=2}^{k-1}\frac{\#([i!,(i+1)!)\cap\mathbb{N})}{k}+\frac{\#([k!,\xi]\cap\mathbb{N})}{k}-1\\ &=\frac{\#([2,\xi]\cap\mathbb{N})}{k}-1\geq\frac{\xi}{2\nu(\xi)},\end{split}$
where $k\geq 1$ is the integer with $\xi\in[k!,(k+1)!)$. This verifies (3.3).
The ordering of elements of $S$ as in the statement of Lemma 3.2 implies that for any $N\geq 1$ with $S\cap[1,N]\neq\emptyset$, we have
$(S_{*}\cap[1,N])=\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}.$
By the second inequality in (3.3), for all sufficiently large $N\geq 1$ we have
$\frac{\#(S_{*}\cap[1,N])}{\#(S\cap[1,N])}=\frac{\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}}{\#(S\cap[1,N])}\leq\frac{3}{\nu(\#(S\cap[1,N]))}.$
@@
3.3. Proof of Proposition 3.1 (Step 2)
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha\geq 1$ and write $S=\{a_{n}\colon n\in\mathbb{N}\}$, $a_{1}<a_{2}<\cdots.$ Let $S_{*}$ be a subset of $S$ for which the conclusion of Lemma 3.2 holds. It suffices to show that $\dim_{\rm H}R_{t,L}(S_{*})\geq 1/(2\alpha)$ for some $t,L$ with $t>L>1$ and $t>2$.
By the ordering of elements of $S$ as above, for any $N\geq 1$ with $S\cap[1,N]\neq\emptyset$ we have $(S_{*}\cap[1,N])=\#\{1\leq n\leq\#(S\cap[1,N])\colon a_{n}\in S_{*}\}$. Moreover (3.3) gives
$\begin{split}\frac{\#(S\cap[1,N])}{2\nu(\#(S\cap[1,N]))}\leq\#(S_{*}\cap[1,N])&\leq\frac{3\#(S\cap[1,N])}{\nu(\#(S\cap[1,N]))}.\end{split}$ (3.4)
@@
Using (3.4), (3.6) and then (3.7), (3.8), for every $n\in\mathbb{N}$ we have
$\begin{split}\#(S_{*}\cap[t^{n},Lt^{n}])&\geq\#(S_{*}\cap[1,Lt^{n}])-\#(S_{*}\cap[1,t^{n}])\\ &\geq\frac{\#(S\cap[1,Lt^{n}])}{2\nu(\#(S\cap[1,Lt^{n}]))}-\frac{3\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}\\ &\geq\frac{C_{1}C_{2}^{-1}2^{-\beta}L^{1/\alpha}\#(S\cap[1,t^{n}])}{2(\nu(\#(S\cap[1,t^{n}])+1)}-\frac{3\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}.\end{split}$
@@
$(S\cap[1,t^{n}])\left(\frac{C_{1}C_{2}^{-1}2^{-\beta}L^{1/\alpha}}{3\nu(\#(S\cap[1,t^{n}]))}-\frac{3}{\nu(\#(S\cap[1,t^{n}]))}\right),$
which equals $(S\cap[1,t^{n}])/\nu(\#(S\cap[1,t^{n}]))$ by (3.5). Set $\rho=\beta+\lambda$. Using (3.9) and the assumption that $S$ has polynomial density, we obtain
$\begin{split}\#(S_{*}\cap[t^{n},Lt^{n}])\geq\frac{\#(S\cap[1,t^{n}])}{\nu(\#(S\cap[1,t^{n}]))}&\geq C_{1}\frac{t^{\frac{n}{\alpha}}}{(n\log t)^{\beta}}\cdot\frac{1}{C_{1}(n\log t)^{\lambda}}\\ &=\frac{t^{\frac{n}{\alpha}}}{(n\log t)^{\rho}}.\end{split}$ (3.10)
@@
where the bar denotes the closure. Since $t>L$, the intervals $t^{n},Lt^{n}]$, $n\in\mathbb{N}$ are pairwise disjoint. The Moran fractal $F$ is contained in $R_{t,L}(S_{*})$ up to countable sets consisting of the boundary points of the level intervals. For every $n\geq 2$ and every (n1)(n-1)-th level interval of $F$, the number of $n$-th level intervals contained in that interval is at least $\lfloor t^{\frac{n}{\alpha}}/(n\log t)^{\rho}\rfloor$} by (3.10). Since $S_{*}$ does not contain consecutive integers, these $n$-th level intervals are separated at least by some $n$-th fundamental intervals. Hence, the length of each gap between the $n$-th level intervals is at least
$\min\{|I(a_{1},\ldots,a_{n})|\colon t^{i}\leq a_{i}\leq Lt^{i}\ \text{ for }1\leq i\leq n\}.$
By Lemma 2.1, $t>L$ and $t>2$, we have
@@
So, $F$ is a Moran fractal with parameters $({\lfloor t^{\frac{n}{\alpha}}/(n\log t)^{\rho}\rfloor})_{n=1}^{\infty}$, $((1/2)\prod_{i=1}^{n}(Lt^{i+1})^{-2})_{n=1}^{\infty}$. Lemma 2.2 gives
@@
For two eventually positive functions $f$, $g$ on $\mathbb{N}$, let us use the expression $f(n)\sim g(n)$ to indicate that $f(n)/g(n)\to 1$ as $n\to\infty$. Direct calculations show that
$\begin{split}\log\left(2\prod_{i=1}^{n-1}{\left\lfloor\frac{t^{\frac{i}{\alpha}}}{(i\log t)^{\rho}}\right\rfloor}\right)&=\log 2+\sum_{i=1}^{n-1}\log\left\lfloor\frac{t^{\frac{i}{\alpha}}}{(i\log t)^{\rho}}\right\rfloor\\ &\sim\log 2+\sum_{i=1}^{n-1}\left(\log t^{\frac{i}{\alpha}}-\log(i\log t)^{\rho}\right)\sim\frac{1}{2\alpha}n^{2}\log t,\end{split}$
@@
In [35], Ramharter proved $\dim_{\rm H}J=\dim_{\rm H}E=1/2$. Combining Proposition 3.1 with statements on convergence exponent in Section 2.3, we obtain a refinement of Ramharter’s result.
Proposition 3.4.
If $S\subset\mathbb{N}$ has polynomial density with exponent $\alpha\geq 1$, then
$\begin{split}\dim_{\rm H}\{x\in J\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}&=\dim_{\rm H}\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}\\ &=\frac{1}{2\alpha}.\end{split}$
@@
3.4. Adding digits into seed sets (Step 3)